the sum of 2 integers is 25. One less than twice the larger integer is equal to one more than four times the smaller integer. Find the integers

Answer by CubeyThePenguin(3113)

x = larger integer
y = smaller integer

x + y = 25
2x - 1 = 4y + 1

Rearranging the first equation, we get x = 25 - y. We can substitute this into the second equation.

2x - 1 = 4y + 1
2(25 - y) - 1 = 4y + 1
50 - 2y - 1 = 4y + 1
48 = 6y
8 = y

The integers are 17 and 8.